# Free energy barrier and thermal-quantum behavior of sliding bilayer graphene

### v3

@@ I Introduction
There has been intense research on graphene since its discovery in 2004, and also on multilayer graphene systems. Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)Koshino (2010), and superconductivity has been observed in RG in 2021Zhou et al. (2021). Several works recently studied, both experimentally and theoretically, under which conditions RG can form Yankowitz et al. (2014); Geisenhof et al. (2019); Bouhafs et al. (2021); Gao et al. (2020); Li et al. (2020); Kerelsky et al. (2021). In particular, some of us unveiled how the stacking of graphene layers changes under shear stress Nery et al. (2020), and obtained that it produces many consecutive layers of RG. In fact, all samples with long-range RG that have been observed so far have been subject to some form of shear stress Lin et al. (2012); Henni et al. (2016); Henck et al. (2018); Yang et al. (2019); Shi et al. (2020); Bouhafs et al. (2021).
@@
To move graphene layers with respect to each other, they have to overcome a free energy barrier that separates different minima. When doing so, they experiment “stick-slip” motion: they move gradually with increasing shear stress, followed by sudden rearrangements and decrease of the stressLiu et al. (2012). The maximum shear-stress to “unlock” layers is correlated with the height of the barrier. Such barrier corresponds to the activation energy of a metastable state, where a graphene flake sitting on top of a large graphene layer is displaced relative to the equilibrium position.
@@
Superlubricity is a another phenomenon where the energy landscape corresponding to the relative displacement of layers is also central. In this mechanism, which has attracted a lot attention(Liu et al., 2014; Sinclair et al., 2018), particularly in carbon-based systems like multilayer graphene and graphite, incommensurate surfaces (i.e. atomic lattices without matching periodicity) lead to extremely low friction. Yet another phenomena involving sliding surfaces is thermolubricity, which refers to the assistance from thermal excitations to overcome barriers. This term has been coined by some works that have studied the transition from stick-slip motion into stochastic fluctuations in friction at sufficiently high temperaturesBaykara et al. (2018). Despite the interest in sliding surfaces, there are no studies that account for both temperature and quantum effects of free-energy barriers in macroscopic systems. A few works have studied quantum effects in free-energy barriers, including zero-point renormalization and tunneling, but of systems involving hydrogen molecules, including dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al. (2012); McIntosh et al. (2013); Cendagorta et al. (2016).
@@
The variation of the potential along a periodic crystal surface, combined with the attraction between layers, prevents free sliding between them. This phenomenon has been extensively studied in various systems, like graphiteDienwiebel et al. (2004), MoS2Li et al. (2016), silicon on goldLiu et al. (2015), among others. When using a tip with the goal of sliding layers, normal forces need not be large. For pressures much lower than 1GPa, the alteration of the free-energy landscape is minimal (see SI of Ref. Nery et al., 2020), so the relaxed barrier remains the dominant contribution that impedes sliding. In addition, some works on superlubricity have observed that friction has an extremely weak dependence with the load, attributed for exampleDienwiebel et al. (2004) to a contact area that remains nearly constant in the force range applied. Here, we focus on calculating the height of the barrier, which should be approximately proportional to the force –given by the maximum slope of the free energy profile– that is needed for a layer to move from one minima to another in commensurate systems.
@@
A quantity closely related to the free energy barrier is the shear frequency, corresponding to the mode in which layers move rigidly in-plane and out of phase around the equilibrium position. The free energy barrier separates two equivalent minima, so the barrier and shear frequency are expected to have a similar temperature dependence (a lower curvature at the minima corresponds to a lower shear frequency and a lower free energy, at least close to the minima). The value of the shear frequency in graphite is a relatively well-established quantity, with experimental valuesLebedeva et al. (2017) oscillating between 42 - 45 cm-1, using for example inelastic X-ray scatteringMohr et al. (2007), neutron-coherent inelastic scatteringNicklow et al. (1972), Raman spectraTan et al. (2012); Hanfland et al. (1989), and coherent phonon spectroscopy Boschetto et al. (2013). In bilayer graphene, there are some experiments with values around 30 cm-1; for example, 28 $\pm$ 3 cm-1 in Refs. S Amelinckx, 1965; Boschetto et al., 2013, and 32 cm-1 in Ref. Tan et al., 2012. The temperature dependence of the shear modes, however, has been barely studied. Few experiments measure the temperature dependence of the shear mode, using Raman scattering in folded 2+2 (4 layer) grapheneCong and Yu (2014), or femtosecond pump-probe spectroscopy in graphiteMishina et al. (2000).
@@
Another aspect of multilayer graphene and graphite that still needs further investigation and that serves as a consistency check on the adequacy of our approach is thermal expansion (TE). Ref. McQuade et al., 2021 measured it in bilayer and trilayer graphene for the first time. Ref. Mittal et al., 2021 uses the quasi-harmonic approximation (QHA), which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972). Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al. (2018). However, in-plane PIMD results of Ref. Herrero and Ramírez, 2020 vastly overestimate the TE.
@@
Motivated by the aforementioned phenomena, especially by the possibility of facilitating sliding through temperature, we focus on thermal effects on the free-energy barrier. This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position. Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Errea et al. (2014); Monacelli et al. (2021). First, in Sec. II, we briefly introduce SSCHA and the QHA, and explain how the barrier is determined. Then, in Sec. III, we present our main results and compare to the height of the barrier with fixed atoms. Subsequently, using also the QHA, we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available. Finally, we determine the in-plane and out-of-plane TE coefficient in bilayer graphene and graphite, and compare our results to previous theoretical and experimental work. In Sec. IV, we summarize the conclusions of our work.
@@ II Theoretical and Computational Framework
Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier. A more detailed description of SSCHA can be found in Refs. Errea et al., 2014 or Monacelli et al., 2021.
@@ II.1 Quasi-harmonic approximation (QHA)
$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N_{q}}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N_{q}}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right).\\ \end{split}$ (1)
@@
The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy. Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion. Using first-principles calculations, phonons can be determined using density functional perturbation theory, which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells. However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature. For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter. Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see the Supplementary Information (SI), Fig. S11), then an acoustic mode becomes unstable and the QHA is ill-defined. These issues are not present in SSCHA.
@@ II.2 Stochastic self-consistent harmonic approximation (SSCHA)
Due to a variational principle, $F_{H}\leq\mathcal{F}_{H}(\mathcal{H})$. The parameters of the harmonic Hamiltonian $\mathcal{H}$ include the centroids of the atoms, which correspond to the lattice parameters $a_{i}$ of the QHA, but it also includes the interatomic force constant matrix, which contains the information of all the phonon frequencies and accounts for the true anharmonicities. Thus, SSCHA has additional parameters, and is expected in general to give better results than QHA. Also, QHA does not satisfy a variational principle. It can be shown that
[…]
@@
The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy. In a purely harmonic system, such frequencies correspond to the physical frequencies. But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al. (2021), which we do in Sec. III.2 to obtain the temperature dependence of the shear frequency. Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the SI, Fig. S6.
@@ III.1 Barrier
We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref. Libbi et al., 2020 and Ref. Wen et al., 2018, respectively. The out-of-plane interactions determine the barrier. The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig. 1, which gives good results when comparing to theory and experiment (see Table 1).
Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical at $T=0$ K) between the interatomic potential we used in this workLibbi et al. (2020); Wen et al. (2018), LDA, and theoretical calculations and experiments from previous works. The values are a little bit lower but close to that of adiabatic-connection fluctuation-dissipation theorem within the random phase approximation (ACFDT-RPA)Zhou et al. (2015) and experimental dataLebedeva et al. (2017).
@@
37 42 42-45 Lebedeva et al. (2017)
The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$. For layers to slide with respect to each other, they have to overcome a barrier $V_{\mathrm{SP}}$. SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable. In order to determine temperature effects on the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP, using the interpolation method described in Sec. II.2. The barrier as a function of temperature can be seen in blue in Fig. 2.
@@
Figure 2: Barrier as a function of temperature. It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig. 1), to about 0.8 meV/atom, corresponding to a large reduction of about 35% due to phonons. The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K. The lines are a guide to the eye, as in the other figures of this work.
@@
The barrier at 100 K is reduced by a substantial 35% with respect to the value at fixed nuclei. We expect intralayer modes to contribute little to the reduction of the barrier, since they should not change much when displacing the layers. On the other hand, modes involving the relative motion of the layers should be much more affected. Indeed, intralayer modes are virtually identical, so they do not contribute to a change in the barrier. Regarding the interlayer modes, the breathing mode changes little, while the shear modes do differ significantly (see Fig. S4). Thus, the shear modes are responsible for the large reduction of the barrier. This makes sense intuitively: modes that involve a relative displacement of the layers should facilitate sliding. In the fixed ion calculation, the system is exactly at SP, but in reality, the system is also exploring surrounding configurations, which have a lower energy.
@@
Remarkably, our results show improved agreement with experiments. In a previous workNery et al. (2020), using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref. Liu et al., 2012) –a benchmark for defect free single crystal graphiteBlomquist et al. (2019)– and 0.1 GPa (Ref. Dienwiebel et al., 2004). Assuming that the maximum shear stress decreases by 35% in line with the barrier, we now obtain 0.13 GPa, in excellent agreement with experimentsDienwiebel et al. (2004); Liu et al. (2012).
[…]
@@
The size of the graphene flakes considered in our work in principle correspond to nanometers up to possibly micrometer scales. In clean samples with micrometer sized flakes, different regions can have different stackings (AB and AC)Henck et al. (2018); Jiang et al. (2018), and a domain wall in the metastable state separates both regions. The displacement of layers in this case is more complex and is not considered in our work. We expect shear stress to be lower in this case, since the metastable state is already present. Defects and stacking faults also lead to lower friction due to incommensurabilities, so the shear stress of 0.14 GPa corresponds to the maximum expected valueBlomquist et al. (2019).

### v4

@@ I Introduction
There has been intense research on graphene since its discovery in 2004, and also on multilayer graphene systems. Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)1, and superconductivity has been observed in RG in 20212. Several works recently studied, both experimentally and theoretically, under which conditions RG can form 3, 4, 5, 6, 7, 8, given that AB-stacked graphene is the more common occurring phase. In particular, some of us unveiled how the stacking of graphene layers changes under shear stress 9, and obtained that it produces many consecutive layers of RG. In fact, all samples with long-range RG that have been observed so far have been subject to some form of shear stress 10, 11, 12, 13, 14, 5.
@@
To move graphene layers with respect to each other, they have to overcome a free energy barrier that separates different minima. When doing so, they experiment “stick-slip” motion: they move gradually with increasing shear stress, followed by sudden rearrangements and decrease of the stress15. The maximum shear-stress to “unlock” layers is correlated with the height of the barrier. Such barrier corresponds to the activation energy of a metastable state, where a graphene flake sitting on top of a large graphene layer is displaced relative to the equilibrium position.
@@
Superlubricity is a another phenomenon where the energy landscape corresponding to the relative displacement of layers is also central. In this mechanism, which has attracted a lot attention16, 17, particularly in carbon-based systems like multilayer graphene and graphite, incommensurate surfaces (i.e. atomic lattices without matching periodicity) lead to extremely low friction. Yet another phenomena involving sliding surfaces is thermolubricity, which refers to the assistance from thermal excitations to overcome barriers. This term has been coined by some works that have studied the transition from stick-slip motion into stochastic fluctuations in friction at sufficiently high temperatures18. Despite the interest in sliding surfaces, there are no studies that account for both temperature and quantum effects of free-energy barriers in macroscopic systems. A few works have studied quantum effects in free-energy barriers, including zero-point renormalization and tunneling, but of systems involving hydrogen molecules, including dissociative adsorption19 and diffusion20, 21, 22.
@@
The variation of the potential along a periodic crystal surface, combined with the attraction between layers, prevents free sliding between them. This phenomenon has been extensively studied in various systems, like graphite23, MoS224, silicon on gold25, among others. When using a tip with the goal of making layers slide, normal forces need not be large. For pressures much lower than 1GPa, the alteration of the free-energy landscape is minimal (see SI of Ref. 9), so the relaxed barrier remains the dominant contribution that impedes sliding. In addition, some works on superlubricity have observed that friction has an extremely weak dependence with the load, attributed for example23 to a contact area that remains nearly constant in the force range applied. Here, we focus on calculating the height of the barrier, which should be approximately proportional to the force –given by the maximum slope of the free energy profile– that is needed for a layer to move from one minima to another in commensurate systems.
@@
A quantity closely related to the free energy barrier is the shear frequency, corresponding to the mode in which layers move rigidly in-plane and out of phase around the equilibrium position. The free energy barrier separates two equivalent minima, so the barrier and shear frequency are expected to have a similar temperature dependence (a lower curvature at the minima corresponds to a lower shear frequency and a lower free energy, at least close to the minima). The value of the shear frequency in graphite is a relatively well-established quantity, with experimental values26 oscillating between 42 - 45 cm-1, using for example inelastic X-ray scattering27, neutron-coherent inelastic scattering28, Raman spectra29, 30, and coherent phonon spectroscopy 31. In bilayer graphene, there are some experiments with values around 30 cm-1; for example, 28 $\pm$ 3 cm-1 in Refs. 32, 31, and 32 cm-1 in Ref. 29. The temperature dependence of the shear modes, however, has been barely studied. Few experiments measure the temperature dependence of the shear mode, using Raman scattering in folded 2+2 (4 layer) graphene33, or femtosecond pump-probe spectroscopy in graphite34.
@@
Another aspect of multilayer graphene and graphite that still needs further investigation and that serves as a consistency check on the adequacy of our approach is thermal expansion (TE). Ref. 35 measured it in bilayer and trilayer graphene for the first time. Ref. 36 uses the quasi-harmonic approximation (QHA), which underestimates the out-of-plane TE at low temperatures37 and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperatures37 and overestimates it at high temperatures38. Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)39, but while the TE flattens at 500 K in graphite, experimental values continue increasing38, 40. However, in-plane PIMD results of Ref. 39 vastly overestimate the TE.
@@
Motivated by the aforementioned phenomena, especially by the possibility of facilitating sliding through temperature, we focus on thermal effects on the free-energy barrier. This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position. Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)41, 42. First, in Sec. II, we briefly introduce SSCHA and the QHA, and explain how the barrier is determined. Then, in Sec. III, we present our main results and compare to the height of the barrier with fixed atoms. Subsequently, using also the QHA, we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available. Finally, we determine the in-plane and out-of-plane TE coefficient in bilayer graphene and graphite, and compare our results to previous theoretical and experimental work. In Sec. IV, we summarize the conclusions of our work.
@@ II Theoretical and Computational Framework
Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier. A more detailed description of SSCHA can be found in Refs. 41 or 42.
@@ II.1 Quasi-harmonic approximation (QHA)
[…]
@@
The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy. Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion. Using first-principles calculations, phonons can be determined using density functional perturbation theory, which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells. However, it does not account for the “true anharmonicity”43, present at fixed values of the parameters when varying temperature. For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter. Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see the Supplementary Information (SI), Fig. S12), then an acoustic mode becomes unstable and the QHA is ill-defined. These issues are not present in SSCHA.
@@ II.2 Stochastic self-consistent harmonic approximation (SSCHA)
Due to a variational principle, $F_{H}\leq\mathcal{F}_{H}(\mathcal{H})$. The parameters of the harmonic Hamiltonian $\mathcal{H}$ include the centroids of the atoms (their average positions), which correspond to the lattice parameters $a_{i}$ of the QHA, but it also includes the interatomic force constant matrix, which contains the information of all the phonon frequencies and accounts for the true anharmonicities. Thus, SSCHA has additional parameters, and is expected in general to give better results than QHA. Also, QHA does not satisfy a variational principle. It can be shown that
$\mathcal{F}_{H}(\mathcal{H})=F_{\mathcal{H}}+\int d\mathbf{R}[V(\mathbf{R})-\mathcal{V}(\mathbf{R})]\rho_{\mathcal{H}}(\mathbf{R}),$ (4)
@@
The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy. In a purely harmonic system, such frequencies correspond to the physical frequencies. But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)42, which we do in Sec. III.2 to obtain the temperature dependence of the shear frequency. Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the SI, Fig. S7.
Since the centroids of the atoms are also parameters in SSCHA, from the minimization of the free energy we additionally obtain the lattice parameters as a function of temperature, allowing us to study their thermal expansion. This is discussed in Sec. III C.
@@ III.1 Barrier
We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref. 44 and Ref. 45, respectively. The out-of-plane interactions determine the barrier. The energy profile of a layer moving with respect to the other along the bond direction, shown in Fig. 1, is obtained by fixing the in-plane positions of the layers and relaxing the interlayer distance until the out-of-plane components of the atomic forces vanish, and gives good agreement with theory and experiment (see Table 1).
Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical at $T=0$ K) between the interatomic potential we used in this work44, 45, LDA, and theoretical calculations and experiments from previous works. The values are a little bit lower but close to that of adiabatic-connection fluctuation-dissipation theorem within the random phase approximation (ACFDT-RPA)46 and experimental data26.
@@
37 42 42-45 26
The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$ Å. For layers to slide with respect to each other, they have to overcome a barrier $V_{\mathrm{SP}}$. Is is worth noting that although the profile is for bilayer graphene, some of us previously showed in first-principles calculations that using 6 layers – 3 layers displaced relative to the other 3 – the profile is almost identical (see SI of Ref. 9). Since interactions with subsequent layers are even weaker, the barrier should stay essentially unchanged with additional layers. In addition, in Ref. 47 we obtained that energy differences between different functionals for different stacking sequences, including van-der-Waals corrections, differ by around 0.01 meV/atom. Thus, the barrier in the bulk limit should correspond closely to the bilayer one.
SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable. In order to determine temperature effects on the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP, using the interpolation method described in Sec. II.2. The barrier as a function of temperature can be seen in blue in Fig. 2.
@@
Figure 2: Barrier as a function of temperature. It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig. 1), to about 0.8 meV/atom, corresponding to a large reduction of about 35% due to phonons. The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K. The lines are a guide to the eye, as in the other figures of this work. The standard quantum SSCHA result (blue) and its classical limit (red) give almost identical results, showing quantum effects are negligible above 100 K.
@@
The barrier at 100 K is reduced by a substantial 35% with respect to the value at fixed nuclei. We expect intralayer modes to contribute little to the reduction of the barrier, since they should not change much when displacing the layers. On the other hand, modes involving the relative motion of the layers should be much more affected. Indeed, intralayer modes are virtually identical, so they do not contribute to a change in the barrier. Regarding the interlayer modes, the breathing mode changes little, while the shear modes do differ significantly (see Fig. S4). Thus, the shear modes are responsible for the large reduction of the barrier. This makes sense intuitively: modes that involve a relative displacement of the layers should facilitate sliding. In the fixed ion calculation, the system is exactly at SP, but in reality, the system is also exploring surrounding configurations, which have a lower energy. In an analogous fashion, there is a barrier in real space, since the interlayer distance at AB is lower relative to SP. At 100 K, this barrier gets similarly reduced by 38%, from $0.029\,\text{\AA }$ to $0.018\,\text{\AA }$, and also shows a temperature dependence similar to that of the free energy barrier (see Figs. S5 and S6).
@@
Remarkably, our results show improved agreement with experiments. In a previous work9, using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref. 15) –a benchmark for defect free single crystal graphite48– and 0.1 GPa (Ref. 23). Assuming that the maximum shear stress decreases by 35% in line with the barrier, we now obtain 0.13 GPa, in excellent agreement with experiments23, 15.
Since 100 K – about 70 cm-1 in wavenumber units – is a relatively low temperature, quantum fluctuations may play a role. To assess their impact on the free energy barrier, we also performed calculations with the classical distribution, corresponding to the classical limit $\hbar\omega/kT\rightarrow 0$. In this way, $2n_{s}+1\rightarrow 2T/\omega$, and the quantum (Bose-Einstein) probability distribution for ionic configurations reduces to the classical Boltzmann distribution (an explicit expression for the distribution is included in the SI). We can see the SSCHA results in the classical limit in red in Fig 2, which are virtually identical to the SSCHA quantum calculations, showing that quantum effects are negligible above $100\text{\,}\mathrm{K}$. The reason is that shear modes responsible for the barrier reduction have very low energies, and even a modest temperature can populate such modes. This is not a general feature, as other systems present significant differences between classical and quantum free energy barriers, as in materials formed by light hydrogen atoms. Such quantum effects have been observed at 100 K (Ref. 19), 250 K (Ref. 21), and even higher temperatures20.
@@
The size of the graphene flakes considered in our work in principle correspond to nanometers up to possibly micrometer scales. In clean samples with micrometer sized flakes, different regions can have different stackings (AB and AC)12, 49, and a domain wall in the metastable state separates both regions. The displacement of layers in this case is more complex and is not considered in our work. We expect shear stress to be lower in this case, since the metastable state is already present. Defects and stacking faults also lead to lower friction due to incommensurabilities, so the shear stress of 0.14 GPa corresponds to the maximum expected value 48.
When it comes to displacing many layers in multilayer graphene or graphite, the change of the barrier could in principle differ, since it depends on the phonon frequencies (and interlayer modes depend noticeably on the number of layers). For example, graphite exhibits a larger shear frequency at equilibrium, since a layer experiences primarily the restoring force of two layers instead of one (see the next subsection for more details). However, layers are typically not simultaneously in unstable positions relative to both nearest layers, so, as noted earlier, the barrier remains approximately unchanged in bulk. Since the frequency at AB increases while remaining similar at SP, the barrier reduction in bulk might be even greater than in the bilayer case. Other variations may arise from the choice of the interlayer interatomic potential: as shown in Table I, the RPA barrier is higher, indicating that the curvature around SP could be larger. Thus, vibrations around SP with more accurate interlayer potentials might yield a larger absolute barrier reduction, although the relative change may still be similar.